# A New Method to Simulate Dark Matter–Baryon Interactions and Application to an Isolated Disk Galaxy

### 25 Mar 2026 v1

@@ Appendix C Equivalence of the two approaches
$=\frac{m_{\chi}}{M_{i}}\,\int{\rm d}^{3}x\,{\rm d}^{3}v\,\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t\,\bm{v}$ (C2)
@@
$=M_{i}\,\sum_{j}\frac{M_{j}}{m_{\chi}+\mu_{j}}\,g_{ij}\,\bigg[\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,\mathcal{B}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})-\bm{v}_{i}\cdot(\bm{v}_{i}-\bm{V}_{j})\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\bigg]\,\Delta t$ (C11)
@@
As we scale up to many dark matter and many gas particles, these relationships continue to hold for each individual pair. This is because both the full expectation values we have just computed and the baryonic rates-of-change simply sum over neighbors of the other type. Hence, for each pair $(i,j)$, equal and opposite terms appear in the expectation values for the change to $i$ and in the rates-of-change of $j$.
@@ Appendix D Computation of overlap integral
For a given softening kernel, the integral which remains can be tabulated in terms of two parameters: $\delta$ and $h_{j}/h_{i}$.
@@
The range of $\delta$ can also be considered. We assume $W$ is zero beyond the smoothing length, so the overlap is zero if $d>h_{i}+h_{j}$. Thus, the range of $\delta$ which gives non-zero overlap can be found:
@@
Appendix E Test details
@@
We perform this test for many choices of these quantities, detailed in Table 2. The standard setup we choose is $\rho_{B}=\rho_{\chi}=10^{10}\ M_{\odot}\,{\rm kpc}^{-3}$, $T_{B}^{0}=10^{4}\ {\rm K}$, $T_{\chi}^{0}=10^{6}\ {\rm K}$, and $V_{\chi}^{0}=200\ {\rm km}\,{\rm s}^{-1}$, which is similar to the conditions near the center of our isolated disk galaxy. The simulations are then analyzed by computing the total momentum (along $x$), the temperature, and the energy of each component for each snapshot.
Name $m_{\chi}$ $\sigma_{0}$ $n$ $V_{\chi}^{0}$ $T_{\chi}^{0}$ $T_{B}^{0}$ $\rho_{B}$ $\rho_{\chi}$ $N_{B}$ $N_{\chi}$
${\rm cm}^{2}$ ${\rm km}\,{\rm s}^{-1}$ K $M_{\odot}\,{\rm kpc}^{-3}$
@@
F25 §3.4 $\bm{m_{p}}$ $\bm{1.6726\times 10^{-23}}$ 0 $\bm{0}$ $\bm{72.69}$ $\bm{484.6}$ $\bm{10^{7}}$ $\bm{10^{7}}$ $\bm{10^{5}}$ $\bm{21^{3}}$
Table 2: Parameter choices for tests; those which differ from the “fiducial” test are indicated in bold.
@@
$\frac{{\rm d}V_{\chi}}{{\rm d}t}$ $=\frac{\rho_{B}}{m_{\chi}+m_{B}}\,\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,(V_{B}-V_{\chi}),$ (E1)
$\frac{{\rm d}V_{B}}{{\rm d}t}$ $=\frac{\rho_{\chi}}{m_{\chi}+m_{B}}\,\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,(V_{\chi}-V_{B}),$ (E2)
$\frac{{\rm d}T_{\chi}}{{\rm d}t}$ $=\frac{\rho_{B}}{(m_{\chi}+m_{B})\,v_{\rm th}^{2}}\left[\mathcal{B}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{B}-T_{\chi}}{m_{\chi}+m_{B}}+\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{\chi}}{m_{\chi}}\,V_{\chi B}^{2}\right],$ (E3)
$\frac{{\rm d}T_{B}}{{\rm d}t}$ $=\frac{\rho_{\chi}}{(m_{\chi}+m_{B})\,v_{\rm th}^{2}}\left[\mathcal{B}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{\chi}-T_{B}}{m_{\chi}+m_{B}}+\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{B}}{m_{B}}\,V_{\chi B}^{2}\right],$ (E4)
@@
${\rm JB}=\frac{1}{6}\,\bigg[S^{2}+\frac{1}{4}\,\big(K-3\big)^{2}\bigg],$ (E5)
where $S$ is the skewness and $K$ is the kurtosis of the velocities. (Note that typically this is multiplied by the number of particles, which we forego in order to compare simulations at different resolutions.) Large values of $\rm JB$} mean large departures from Maxwell-Boltzmann. For the analysis of our simulations, we compute $\rm JB$} for the dark matter velocities independently along $x$, $y$, and $z$ for each snapshot. We then take the maximum value seen at any time along any axis as our guide for whether the Maxwell-Boltzmann assumption is satisfactory for a given test.
@@
The simulation results and naïve Maxwell-Boltzmann predictions are summarized in Figures 10 through 12, which show the evolution of the momenta (left panel), temperatures (middle), and energies (right). Each panel shows the simulation results in blue (dark matter), pink (gas), and grey (total) solid lines, the naïve M-B prediction in a dotted green line, and the initial value of the total momentum or energy as a black dashed line. The maximum value of the Jarque-Bera statistic during the simulation is also reported for each test. Figure 10 shows the “fiducial” simulation, run with ten different random seeds in both the initial conditions generator and the seed inside GIZMO. These help to give an idea of the scatter in the results due to the stochasticity of our method. Figure 11 shows the tests which vary the experimental setup (velocities, temperature, densities, resolutions), all using the same dark matter model: mass $m_{\chi}=2\,m_{p}$, velocity-independent cross section $\sigma=10\ {\rm mb}$. Figure 12 shows the remaining tests, which vary the dark matter model but all use the same experimental setup (note that we change the initial dark matter temperature to keep the velocity dispersion fixed).
In all simulations except the reduced dark matter resolution ($N_{\chi}=24^{3}$), we see excellent conservation of the global momentum and energy, and the lack of perfect conservation in this one simulation is consistent with increased stochasticity because of the reduced resolution. All simulations show sensible behavior, with the dark matter and baryons equilibrating in both momentum and temperature. In all cases where the dark matter remains relatively Maxwell-Boltzmann (empirically, when $\text{max JB}\lesssim 0.01$), we see excellent agreement between the simulation and the M-B prediction. For tests that attain larger JB values, the M-B prediction deviates from the simulation, usually predicting faster energy and momentum transfer than is observed in the simulation. We stress that this is a failure of the naïve Maxwell-Boltzmann prediction and not of the simulation: for these test setups, the dark matter really does deviate from a Maxwell-Boltzmann distribution. Importantly, this suggests that assuming a Maxwell-Boltzmann distribution can be misleading for situations which arise in the evolution of galaxies. This has been similarly pointed out for the effects of dark matter–baryon interactions on linear cosmology observables (Ali-Haïmoud, 2019).

### 17 Apr 2026 v2

@@ Appendix C Equivalence of the two approaches
$=\frac{m_{\chi}}{M_{i}}\,\int{\rm d}^{3}x\,{\rm d}^{3}v\,\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t\,\bm{v}.$ (C2)
@@
$=M_{i}\,\sum_{j}\frac{M_{j}}{m_{\chi}+\mu_{j}}\,g_{ij}\,\bigg[\frac{\mu_{j}}{m_{\chi}+\mu_{j}}\,\mathcal{B}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})-\bm{v}_{i}\cdot(\bm{v}_{i}-\bm{V}_{j})\,\mathcal{A}(\bm{v}_{i}-\bm{V}_{j};T_{j}/\mu_{j})\bigg]\,\Delta t.$ (C11)
@@
As we extend to many dark matter and many gas particles, these relationships continue to hold for each individual pair. This is because both the full expectation values we have just computed and the baryonic rates-of-change simply sum over neighbors of the other type. Hence, for each pair $(i,j)$, equal and opposite terms appear in the expectation values for the change to $i$ and in the rates-of-change of $j$.
@@ Appendix D Computation of overlap integral
For a given softening kernel, the integral that remains can be tabulated in terms of two parameters: $\delta$ and $h_{j}/h_{i}$.
@@
The range of $\delta$ can also be considered. We assume $W$ is zero beyond the smoothing length, so the overlap is zero if $d>h_{i}+h_{j}$. Thus, the range of $\delta$ that gives non-zero overlap can be found:
@@
Appendix E Regime of validity of the F25 method
As discussed in Section II.5, we have performed a numerical experiment to quantify the regions of validity of the “rare, large-angle scattering” method of Fischer et al. (2025, F25). Here we describe that experiment.
We consider two macroparticles, one dark matter and one baryon, with velocities like those of our test method (see Section III or the more detailed discussion below). For this single pair of macroparticles, assumed to be overlapping with some fiducial value for $g_{ij}$, we perform 100 000 scatterings. Each scattering involves randomly sampling a dark matter velocity from a Maxwell-Boltzmann distribution, a “virtual baryon” velocity from a Maxwell-Boltzmann distribution, and a scattering angle. We then compute the change to the bulk velocity $\bm{V}_{B}$ and the internal specific energy $U_{B}$ of the baryon according to the algorithm described in F25.
We are primarily interested in the energy transfer. We define the total specific energy of the baryon to be
$\mathcal{E}_{B}=\frac{1}{2}\bm{V}_{B}^{2}+U_{B}.$ (E1)
In each trial $i$, the scattering occurs with some probability $P_{i}$, which depends on the sampled velocities. So, the expected change in the total energy of the baryon for trial $i$ is
$\Delta\mathcal{E}_{i}=P_{i}\,(\mathcal{E}_{B}^{\prime}-\mathcal{E}_{B})+(1-P_{i})\,0=P_{i}\,(\mathcal{E}_{B}^{\prime}-\mathcal{E}_{B}).$ (E2)
By averaging this quantity over many trials, we compute the expectation value of the change in energy, integrating by Monte Carlo over scattering angle and the dark matter and baryon velocity distributions:
$\mathbb{E}[\Delta\mathcal{E}_{B}]=\frac{1}{N_{\text{trials}}}\,\sum_{i\in\text{trials}}P_{i}\,(\mathcal{E}_{B}^{\prime}-\mathcal{E}_{B}).$ (E3)
In the F25 algorithm, a scattering is rejected if it results in negative internal energy $U_{B}^{\prime}$. Hence, given all these trial scatterings, the expected change in energy that will actually result from the F25 algorithm is given by averaging over only those trials where $U_{B}^{\prime}>0$:
$\mathbb{E}[\Delta\mathcal{E}_{B}]_{\text{actual}}=\frac{1}{N_{\text{trials}|U_{B}^{\prime}>0}}\,\sum_{i\in\text{trials}|U_{B}^{\prime}>0}P_{i}\,(\mathcal{E}_{B}^{\prime}-\mathcal{E}_{B}).$ (E4)
We can also compute the expected rejection rate, which is the probability-weighted fraction of scatterings that were rejected:
$f_{\text{reject}}=\frac{\sum_{i\in\text{trials}}P_{i}\,\mathbf{1}(U_{B}^{\prime}\leq 0)}{\sum_{i\in\text{trials}}P_{i}},$ (E5)
where $\mathbf{1}(\text{cond})$ is the condition function: 1 if the condition is true and 0 otherwise.
We expect the rejection rate to depend primarily on two mass ratios: that of the macroparticles $M_{\chi}/M_{B}$ and of the physical particles $m_{\chi}/m_{B}$. We therefore conduct this experiment for a wide range of both ratios. The resulting rejection rates and energy change bias fractions (actual divided by expected) are shown as heatmaps in Figure 11. Also shown on these figures is the expected “safe zone” described in F25:
$\frac{M_{\chi}}{M_{B}}<\frac{3}{28}\,\frac{m_{\chi}}{m_{B}}.$ (E6)
Figure 11: The expected rejection rate (left) and energy transfer bias ratio (right) of the Fischer et al. (2025) algorithm as a function of the mass ratios of the true particles, $m_{\chi}/m_{B}$, and of the macroparticles, $M_{\chi}/M_{B}$. The accuracy of the method is highly degraded whenever $M_{\chi}/M_{B}\gtrsim m_{\chi}/m_{B}$. An analytically predicted “safe zone” (Eq. 12, Fischer et al., 2025) is shown hatched with circles.
We find that the rejection rate is zero and the bias fraction is one everywhere in the expected safe zone, and even for $M_{\chi}/M_{B}$ slightly greater than the safe zone threshold. However, performance degrades steeply once $M_{\chi}/M_{B}\gtrsim m_{\chi}/m_{B}$, with the rejection rate becoming 100% almost everywhere. In general, the effect of the rejections is to reduce the number of scatters and therefore to bias the energy transfer downward, which is borne out by the observed bias ratio. In fact, for most of the mass ratios tested, the bias ratio becomes exactly zero, with energy transfer completely shut off. However, there is a small region where the rejections actually enhance the energy transfer; this is because the distribution of energy transfers is broad when the macroparticle mass ratio is similar to that of the physical particles, and the rejection step only cuts off the negative tail of the distribution.
Appendix F Test details
@@
We perform this test for many choices of these quantities, detailed in Table 2. The standard setup we choose is $\rho_{B}=\rho_{\chi}=10^{10}\ M_{\odot}/{\rm kpc}^{3}$, $T_{B}^{0}=10^{4}\ {\rm K}$, $T_{\chi}^{0}=10^{6}\ {\rm K}$, and $V_{\chi}^{0}=200\ {\rm km}/{\rm s}$, which is similar to the conditions near the center of our isolated disk galaxy. The simulations are then analyzed by computing the total momentum (along $x$), the temperature, and the energy of each component for each snapshot.
Name $m_{\chi}$ $\sigma_{0}$ $n$ $V_{\chi}^{0}$ $T_{\chi}^{0}$ $T_{B}^{0}$ $\rho_{\chi}$ $\rho_{B}$ $N_{\chi}$ $N_{B}$
${\rm cm}^{2}$ ${\rm km}/{\rm s}$ K $M_{\odot}/{\rm kpc}^{3}$
@@
F25 §3.4 $\bm{m_{p}}$ $\bm{1.6726\times 10^{-23}}$ 0 $\bm{0}$ $\bm{484.6}$ $\bm{72.69}$ $\bm{10^{7}}$ $\bm{10^{7}}$ $\bm{10^{5}}$ $\bm{21^{3}}$
Table 2: Parameter choices for tests; those that differ from the “fiducial” test are indicated in bold.
@@
$\frac{{\rm d}V_{\chi}}{{\rm d}t}$ $=\frac{\rho_{B}}{m_{\chi}+m_{B}}\,\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,(V_{B}-V_{\chi}),$ (F1)
$\frac{{\rm d}V_{B}}{{\rm d}t}$ $=\frac{\rho_{\chi}}{m_{\chi}+m_{B}}\,\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,(V_{\chi}-V_{B}),$ (F2)
$\frac{{\rm d}T_{\chi}}{{\rm d}t}$ $=\frac{\rho_{B}}{(m_{\chi}+m_{B})\,v_{\rm th}^{2}}\left[\mathcal{B}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{B}-T_{\chi}}{m_{\chi}+m_{B}}+\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{\chi}}{m_{\chi}}\,V_{\chi B}^{2}\right],$ (F3)
$\frac{{\rm d}T_{B}}{{\rm d}t}$ $=\frac{\rho_{\chi}}{(m_{\chi}+m_{B})\,v_{\rm th}^{2}}\left[\mathcal{B}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{\chi}-T_{B}}{m_{\chi}+m_{B}}+\mathcal{A}(V_{\chi B};v_{\rm th}^{2})\,\frac{T_{B}}{m_{B}}\,V_{\chi B}^{2}\right],$ (F4)
@@
${\rm JB}=\frac{1}{6}\,\bigg[S^{2}+\frac{1}{4}\,\big(K-3\big)^{2}\bigg],$ (F5)
where $S$ is the skewness and $K$ is the kurtosis of the velocities. (Note that typically this is multiplied by the number of particles, which we forgo in order to compare simulations at different resolutions.) Large values of $\rm JB$} mean large departures from Maxwell-Boltzmann. For the analysis of our simulations, we compute $\rm JB$} for the dark matter velocities independently along $x$, $y$, and $z$ for each snapshot. We then take the maximum value seen at any time along any axis as our guide for whether the Maxwell-Boltzmann assumption is satisfactory for a given test.
@@
The simulation results and naïve Maxwell-Boltzmann predictions are summarized in Figures 12 through 14, which show the evolution of the momenta (left panel), temperatures (middle), and energies (right). Each panel shows the simulation results in blue (dark matter), pink (gas), and grey (total) solid lines, the naïve M-B prediction in a dotted green line, and the initial value of the total momentum or energy as a black dashed line. The maximum value of the Jarque-Bera statistic during the simulation is also reported for each test. Figure 12 shows the “fiducial” simulation, run with ten different random seeds in both the initial conditions generator and the seed inside GIZMO. These help to give an idea of the scatter in the results due to the stochasticity of our method. Figure 13 shows the tests that vary the experimental setup (velocities, temperature, densities, resolutions), all using the same dark matter model: mass $m_{\chi}=2\,m_{p}$, velocity-independent cross section $\sigma=10\ {\rm mb}$. Figure 14 shows the remaining tests, which vary the dark matter model but all use the same experimental setup (note that we change the initial dark matter temperature to keep the velocity dispersion fixed).
In all simulations except the reduced dark matter resolution ($N_{\chi}=24^{3}$), we see excellent conservation of the global momentum and energy, and the lack of perfect conservation in this one simulation is consistent with increased stochasticity because of the reduced resolution. All simulations show sensible behavior, with the dark matter and baryons equilibrating in both momentum and temperature. In all cases where the dark matter remains relatively Maxwell-Boltzmann (empirically, when $\text{max JB}\lesssim 0.01$), we see excellent agreement between the simulation and the M-B prediction. For tests that attain larger JB values, the M-B prediction deviates from the simulation, usually predicting faster energy and momentum transfer than is observed in the simulation. We stress that this is a failure of the naïve Maxwell-Boltzmann prediction and not of the simulation: for these test setups, the dark matter really does deviate from a Maxwell-Boltzmann distribution. Importantly, this suggests that assuming a Maxwell-Boltzmann distribution can be misleading for situations that arise in the evolution of galaxies. This has been similarly pointed out for the effects of dark matter–baryon interactions on linear cosmology observables (Ali-Haïmoud, 2019).